quadratic equations; completing the sqaure
y^2+2y=2

Answer by ewatrrr(24785)

 
Hi,
completing the square
y^2+2y=2
(y+1)^2 = 2+1
(y+1)^2 = 3
y + 1 = ± sqrt(3)
y = -1 ± sqrt(3) y is -2.732,.732